# An Efficient and Adaptive Watermark Detection System with Tile-based Error Correction

### v1

@@ 4  $s\leftarrow\textsc{SelectTileSize}(i)$;
5  (lat, mem) $\leftarrow\textsc{PredictFromWarmup}(s,\Pi,b_{0},\textsc{Mode})$;
6  push Task(ii, $s$, lat, mem) into $K$;
7
8
@@
10  $S_{p}\leftarrow[\,]$;  load($p$) $\leftarrow 0$;
11
12
@@ 10  $S_{p}\leftarrow[\,]$;  load($p$) $\leftarrow 0$;
14  kappa $\leftarrow$ task in $K$ with max latency;
15  remove kappa from $K$;
@@ 16  $p^{\star}\leftarrow$ stream with minimum load;
17  if load($p^{\star}$)+kappa.lat $\leq(1+\lambda)\cdot\min_{q}$ load($q$) and MemOK($S\cup\{\kappa\},M_{\mathrm{cap}}$) then
18     assign kappa to stream $p^{\star}$;
19     load($p^{\star}$) $\mathrel{+}=$ kappa.lat;
@@
22     (k1, k2) $\leftarrow$ Shard(kappa, $b_{\min}$);
23     assign k1 to stream $p^{\star}$;
24     load($p^{\star}$) $\mathrel{+}=$ k1.lat;
25     push k2 back into $K$;
26
27
28
/* Step 4: Assign uniform mini-batch size */
@@ 30 $m_{\text{unit}}\leftarrow\max(b_{\min},\lfloor B/u\rfloor)$;
31 foreach task kappa in each stream do
32  kappa.mb $\leftarrow m_{\text{unit}}$;
33
34
35return $(S,m_{\text{unit}})$;
Algorithm 2 Resource-aware mini-batch scheduling.
@@
In practice, we first profile stage latencies $(t_{0},t_{1},t_{2})$ and per-sample memory usage during the warm-up phase of our stream allocation strategy proposed in §5.2. Then, for a mini-batch of size $m$ under a baseline batch $b_{0}$ and a stream allocation $(s_{0},s_{1},s_{2})$, we formulate the predicted bottleneck cost as $c(m)=\max\Bigl{\{}\tfrac{t_{0}}{s_{0}},\,\tfrac{t_{1}}{s_{1}},\,\tfrac{t_{2}}{s_{2}}\Bigr{\}}\cdot\tfrac{m}{b_{0}}.$ Scheduling proceeds with a longest-processing-time (LPT) rule: each mini-batch is placed on the stream with the lowest accumulated load, subject to load balance and memory constraints. Low input variance further ensures schedule stability under this rule. If either constraint fails, the mini-batch is split down to a minimum size $b_{\min}$ and rescheduled. In practice, preprocessing is scheduled first, and its outputs feed into downstream queues, so that tiling and decoding automatically inherit the smoothed arrivals established by preprocessing. The global batch $B$ is evenly partitioned into mini-batches (clipped by $b_{\min}$), which amortizes the overhead of frequent launches.
@@
As detailed in Algorithm 2, the scheduler first builds candidate tasks: it initializes an empty pool $K$, then for each image selects a tile size, predicts its latency and memory from warm-up statistics, and pushes a task record into $K$ (Lines2–5). It then initializes streams by creating $P$ empty stream queues $S_{p}$ and zeroing their accumulated loads (Lines6–7). Next, it performs LPT scheduling with balance check: it repeatedly pops the currently most time-consuming task $\kappa$, chooses the stream $p^{\star}$ with the minimum load (Lines9–11), and tests two constraints—load balance with slack $\lambda$ and global memory feasibility (Line12). If both pass, $\kappa$ is assigned to $p^{\star}$ and the stream load is updated by its latency (Lines13–14). Otherwise, $\kappa$ is sharded at the fine granularity of the minimum mini-batch $b_{\min}$ into $(k_{1},k_{2})$; $k_{1}$ is placed onto $p^{\star}$ and $k_{2}$ is returned to the pool for later placement (Lines15–18). The loop continues until all tasks are placed. Finally, the scheduler assigns a uniform mini-batch size: it counts the total number of scheduled tasks across streams, derives $m_{\text{unit}}=\max(b_{\min},\lfloor B/u\rfloor)$ (Line20–21), and sets each task’s mini-batch to $m_{\text{unit}}$ (Line22). It then returns the final stream schedule $S$ and the common mini-batch size $m_{\text{unit}}$ (Line 24).
@@ 7 Evaluation
In this section, we evaluate our QRMark system alongside several popular baseline methods and demonstrate the experimental findings. Our evaluation will be focusing on several aspects: (1) end-to-end system performance comparison among benchmarks, (2) adaptability to the existing watermark detection framework, (3) detection efficiency analysis, and (4) ablation study on the effects of different system design choices. This structure highlights the performance, adaptability, scalability, and expressiveness of QRMark.
@@
Baselines and Metrics: Our benchmarks include two recent and representative works in diffusion model watermarking: Stable Signature [9] and AquaLoRA [8]. Most of QRMark’s optimizations are designed and tailored for Stable Signature. For AquaLoRA, which by default only supports a batch size of one, we minimally patched it to allow variable batch sizes during watermark extraction. To provide a comprehensive evaluation of QRMark’s performance, our metrics include: (i) detection latency and throughput, (ii) detection accuracy without watermark attacks, and (iii) robustness under various adversarial attacks. Additionally, to assess the impact of watermarking on image quality, we report image distortion using the average Peak Signal-to-Noise Ratio (PSNR).
Datasets: During the watermark detection phase, each benchmark is evaluated on the same dataset as used in its original paper. We evaluate Stable Signature based on generations on prompts from the validation set of MS-COCO [19], a large-scale, widely used vision benchmark containing images with rich annotations across 80 object categories. On the other hand, we evaluated AquaLoRA based on generations on selected prompts from partiprompts [45].
@@ 7.1 End-to-End Performance
Experiments on Stable Signature: The comparison of the accuracy of the QRMark and the benchmarks is shown in Table 2. Note that the distortion averages reported in the table correspond to the mean bit accuracy under the set of attacks evaluated in the respective benchmark papers. For Stable Signature, the considered attacks includes: crop, resize, jpeg, brightness, contrast, saturation, sharpness, and overlay_text (full descriptions on those attacks could be found in Table 1). Except for the case of tile size =16=16, where QRMark exhibits relatively lower bit accuracy and average bit accuracy under different distortions, QRMark achieves nearly identical accuracy to the original Stable Signature framework at tile sizes $32$, $48$, $64$, and $80$ and steadily approaches its robustness performance.
In Fig. 6, on average, QRMark achieves a $2.43\times$ throughput speedup over the original Stable Signature pipeline for image watermark detection. As for the end-to-end latency presented in Fig. 7, as batch size increases, the batch latency of both QRMark and Stable Signature grows, but the growth rate of QRMark is significantly slower. At a batch size of 512, the latency of Stable Signature approaches nearly four times that of QRMark, highlighting the superior scalability of QRMark. The improvements in speed and latency are largely attributable to three key algorithmic and system-level optimizations: image tiling, adaptive stream allocation, and resource-aware scheduling. Image tiling reduces both computational and memory costs of decoding, as each tile in our setting is only $1/16$ the size of the original image. To further optimize execution, we employ adaptive stream allocation with resource-aware scheduling, which distributes CUDA streams across pipeline stages in proportion to their measured demands and assigns minibatches for each stream. This ensures that the most time-consuming stages receive more parallel capacity. The resulting multi-stream execution improves GPU utilization, boosts throughput, and lowers latency. Besides, the workload interleaving strategy, kernel fusion (full details in Appendix B.1), and optimizations on RS Correction also contribute to the improvement.
Figure 6: End-to-end image throughput comparison for (a) Stable Signature and (b) AquaLoRA with and without our QRMark under varying image batch sizes on a single NVIDIA H100 GPU.
Figure 7: End-to-end watermark extraction latency comparison for (a) Stable Signature and (b) AquaLoRA with and without our QRMark under varying image batch sizes on a single NVIDIA H100 GPU.
@@
Moreover, as shown in Fig. 6 and  7, QRMark further optimized AquaLoRA’s throughput and latency by fusing the image preprocessing stages into a single CUDA kernel and reducing the intermediate global-memory writes and reads between kernels. QRMark with a batch size of 32 achieves a $2.0\times$ speedup over the original AquaLoRA baseline and a $1.2\times$ speedup when using the same batch size. Notably, when scaling the batch size above 128, we observed that the original AquaLoRA baseline encounters CUDA out-of-memory errors, whereas QRMark continues to sustain high throughput and reduce latency. This further highlights our method’s advantage in large-batch image watermark decoding tasks.
@@ 7.2 Optimization Analysis
Speedup Breakdown and Optimality Analysis: To better illustrate the benefits of individual optimizations and quantify the gap to optimal performance, we present a detailed speedup breakdown and optimality analysis of QRMark in Fig. 8. All experiments use a fixed message length of 48 bits and, when applicable, a tile size of $64$. Enabling tiling provides clear advantages over the Stable Signature baseline: after using large-batch (LB), the baseline improves by at most $1.06\times$, whereas simply augmenting Stable Signature with a tiling strategy already achieves a $1.18\times$ speedup, and reaches up to $1.23\times$ under favorable batch sizes. Finally, adaptive CUDA stream allocation (Allocation) further enhances pipeline concurrency and improves GPU utilization, thereby yielding up to $2.43\times$ end-to-end speedup. These optimizations collectively account for the major observed performance gains.
Figure 8: Optimization breakdown and speedup analysis. (Allocation: Adaptive Streams Allocation, CPU: Reed-Solomon Correction CPU threads Pool, LB: Large Batch, T + F: Tiling + Kernel Fusion)
GPU Utilization: To highlight the efficiency of QRMark in utilizing GPU resources, we profiled average utilization under the default watermark detection setting with a tile size of 64. The Original Stable Signature and its tiled variant exhibited relatively low usage, at only 32.8% and 15.6%, respectively. In contrast, QRMark achieved 66.7% utilization, as its resource-aware CUDA allocation and resource-aware scheduling effectively reduced resource contention and enabled substantially higher GPU occupancy.
7.3 Ablation Study

### v2

@@ 6  $s\leftarrow\textsc{SelectTileSize}(i)$;
7  $(\mathrm{lat},\mathrm{mem})\leftarrow$
8     $\textsc{PredictFromWarmup}(s,\Pi,b_{0},\textsc{Mode})$;
9  push Task(ii, $s$, $\mathrm{lat},\mathrm{mem}$) into $K$;
10
11
@@ 8     $\textsc{PredictFromWarmup}(s,\Pi,b_{0},\textsc{Mode})$;
13  $S_{p}\leftarrow[\,]$;    $\mathrm{load}(p)\leftarrow 0$;
14
15
@@ 13  $S_{p}\leftarrow[\,]$;    $\mathrm{load}(p)\leftarrow 0$;
17  $\kappa\leftarrow$ task in $K$ with max latency;
18  Remove $\kappa$ from $K$;
@@ 19  $p^{\star}\leftarrow$ stream with minimum load;
21 if $\mathrm{load}(p^{\star})+\kappa.\mathrm{lat}\leq(1+\lambda)\cdot\min_{q}\mathrm{load}(q)$
22     and $\textsc{MemOK}(S\cup\{\kappa\},M_{\mathrm{cap}})$ then
23     Assign $\kappa$ to stream $p^{\star}$;
24     $\mathrm{Load}(p^{\star})\mathrel{+}=\kappa.\mathrm{lat}$;
25
@@ 23     Assign $\kappa$ to stream $p^{\star}$;
27     $(k_{1},k_{2})\leftarrow\textsc{Shard}(\kappa,b_{\min})$;
28     Assign $k_{1}$ to stream $p^{\star}$;
29     $\mathrm{Load}(p^{\star})\mathrel{+}=k_{1}.\mathrm{lat}$;
30     Push $k_{2}$ back into $K$;
31
32
33
/* Step 4: Assign uniform micro-batch size */
@@ 35 $m_{\text{unit}}\leftarrow\max(b_{\min},\lfloor B/u\rfloor)$;
36 foreach task $\kappa$ in each stream do
37  $\kappa.\mathrm{mb}\leftarrow m_{\text{unit}}$;
38
39
40return $(S,m_{\text{unit}})$;
Algorithm 2 Resource-aware Micro-batch Schedule.
@@
In practice, we first profile stage latencies $(t_{0},t_{1},t_{2})$ and per-sample memory usage during the warm-up phase of our multi-channel horizontal fusion strategy proposed in §5.2. Then, for a micro-batch of size $m$ under a baseline batch $b_{0}$ and a stream allocation $(s_{0},s_{1},s_{2})$, we formulate the predicted bottleneck cost as $c(m)=\max\Bigl\{\tfrac{t_{0}}{s_{0}},\,\tfrac{t_{1}}{s_{1}},\,\tfrac{t_{2}}{s_{2}}\Bigr\}\cdot\tfrac{m}{b_{0}}.$ Scheduling proceeds with a longest-processing-time (LPT) rule: each micro-batch is placed on the stream with the lowest accumulated load, subject to load balance and memory constraints. Low input variance further ensures schedule stability under this rule. If either constraint fails, the micro-batch is split down to a minimum size $b_{\min}$ and rescheduled. In practice, preprocessing is scheduled first, and its outputs feed into downstream queues, so that tiling and decoding automatically inherit the smoothed arrivals established by preprocessing. The global batch $B$ is evenly partitioned into micro-batches (clipped by $b_{\min}$), which amortizes the overhead of frequent launches.
@@
As detailed in Algorithm 2, the scheduler first builds candidate tasks: it initializes an empty pool $K$, then for each image selects a tile size, predicts its latency and memory from warm-up statistics, and pushes a task record into $K$ (Line 2–5). It then initializes streams by creating $P$ empty stream queues $S_{p}$ and zeroing their accumulated loads (Line 6–7). Next, it performs LPT scheduling with balance check: it repeatedly pops the currently most time-consuming task $\kappa$, chooses the stream $p^{\star}$ with the minimum load (Line 9–13), and tests two constraints—load balance with slack $\lambda$ and global memory feasibility (Line 14). If both pass, $\kappa$ is assigned to $p^{\star}$ and the stream load is updated by its latency (Lines 15–16). Otherwise, $\kappa$ is sharded at the fine granularity of the minimum micro-batch $b_{\min}$ into $(k_{1},k_{2})$; $k_{1}$ is placed onto $p^{\star}$ and $k_{2}$ is returned to the pool for later placement (Line 18–21). The loop continues until all tasks are placed. Finally, the scheduler assigns a uniform micro-batch size: it counts the total number of scheduled tasks across streams, derives $m_{\text{unit}}=\max(b_{\min},\lfloor B/u\rfloor)$ (Line 22–23), and sets each task’s micro-batch to $m_{\text{unit}}$ (Line 24). It then returns the final stream schedule $S$ and the common micro-batch size $m_{\text{unit}}$ (Line 26).
7 Implementation
QRMark is built upon the Stable Signature framework while introducing an abstract and modular architecture that supports both fine-tuning and detection workflows. It adopts a multi-process scheduling design in which detector workers are launched in separate processes, enabling concurrent kernel execution on the GPU. When multiple workers share a single GPU, NVIDIA’s Multi-Process Service (MPS) can be optionally enabled to eliminate excessive context switching and improve intra-GPU parallelism. As is shown in Listing 1, in practice, users can first define a custom watermarking and tiling policy to fit the texture and size of the images, then load the pretrained model and configure key runtime settings such as tiling granularity, the number of GPUs, degree of parallelism, and whether to enable Reed–Solomon decoding. QRMark supports two operating modes. The fine-tuning mode adapts the model to a target dataset and produces images embedded with watermarks. The detection mode verifies and decodes embedded watermarks from input images with efficient tile-based processing.
Listing 1: Custom Watermarking Policy with QRMark.
1import QRMark
2# import other packages... e.g., PyTorch, Triton ...
3# 1. Define Custom Policy Class
4class CustomPolicy(QRMark.Module):
5 def __init__(self, arch="sd-v2.1", bits=48):
6 super().__init__()
7 self.backbone = QRMark.Backbone(arch)
8 self.extractor = QRMark.Extractor(bits)
9 # User-defined strategy example
10 def select_patch(self, x):
11 p = x.unfold(2, 64, 32).unfold(3, 64, 32)
12 idx = p.var((-1,-2)).sum(1).argmax(1)
13 return self.gather(x, idx)
14 # Define the detection forward pass
15 def forward_detect(self, x):
16 x_norm = QRMark.transforms.vqgan_normalize(x)
17 patch = self.select_patch(x_norm)
18 return self.extractor(patch, rs=True)
19# 2. System Configuration
20config = {"num_gpus": 4, "parallel_mode": "data_parallel", "rs": True, ...}
21model = CustomPolicy().to(config)
22# 3. finetune and detect
23model.finetune(datasets, key=secret_key,param**)
24recovered_key = model.detect(suspect_images, param**)
25accuracy = QRMark.verify(recovered_key, secret_key)
@@ 8 Evaluation
In this section, we evaluate our QRMark system alongside several popular baseline methods and demonstrate the experimental findings. Our evaluation will focus on several aspects: (1) end-to-end system performance comparison among benchmarks, (2) adaptability to the existing watermark detection framework, (3) detection efficiency analysis, and (4) ablation study on the effects of different system design choices. This structure highlights the performance, adaptability, scalability, and expressiveness of QRMark.
@@
Baselines and Metrics: Our benchmarks include two recent and representative works in diffusion model watermarking: Stable Signature [fernandez2023stable] and AquaLoRA [feng2024aqualora]. Most of QRMark’s optimizations are designed and tailored for Stable Signature. For AquaLoRA, which by default only supports a batch size of one, we minimally patched it to allow variable batch sizes during watermark extraction. To provide a comprehensive evaluation of QRMark’s performance, our metrics include: (i) detection latency and throughput, (ii) detection accuracy without watermark attacks, and (iii) robustness under various adversarial attacks. Additionally, to assess the impact of watermarking on image quality, we report image distortion using the average Peak Signal-to-Noise Ratio (PSNR).
Datasets: During the watermark detection phase, each benchmark is evaluated on the same dataset used in its original paper. We evaluate Stable Signature using generations from prompts in the MS-COCO validation set [lin2015microsoftcococommonobjects], following its official evaluation protocol. Using only the MS-COCO validation split ensures a fair comparison, as Stable Signature [fernandez2023stable] explicitly reports results on this set. Moreover, evaluating text-to-image models on the MS-COCO validation set is a standard practice in the literature: OpenAI generates samples using captions from the MS-COCO validation set [ramesh2022hierarchical], and Google Brain likewise uses the COCO validation split as the standard benchmark for text-to-image evaluation [saharia2022photorealistic]. We therefore follow this established protocol. AquaLoRA is evaluated using generations from selected prompts in partiprompts [yu2022scalingautoregressivemodelscontentrich], consistent with its original setup.
@@ 8.1 End-to-End Performance
Experiments on Stable Signature: Note that the distortion averages reported in the table correspond to the mean bit accuracy under the set of attacks evaluated in the respective benchmark papers. For Stable Signature, the considered attacks include crop, resize, jpeg, brightness, contrast, saturation, sharpness, and overlay_text (full descriptions of these attacks are provided in Table 1). We adopt this collection because it represents the core set of robustness perturbations used in Stable Signature and related watermarking papers, and therefore provides a representative and comparable evaluation protocol. Importantly, the type of attack does not affect the overall efficiency of the detection pipeline, as all images pass through the same tiling, decoding, and post-processing steps regardless of the applied distortion. Except for the case of tile size $16$, where QRMark exhibits relatively lower bit accuracy, QRMark achieves nearly identical accuracy to the original Stable Signature framework at tile sizes $32$, $48$, $64$, and $80$, and steadily approaches its robustness performance. We use a tile size of $16$ for demonstration and ablation, but such small tiles are uncommon in real deployments; larger tiles (e.g., $64$) provide substantially higher embedding capacity for the 48-bit message and therefore yield near-lossless robustness and better runtime efficiency.
In Fig. 7, on average, QRMark achieves a $2.43\times$ throughput speedup over the original Stable Signature pipeline for image watermark detection. As for the end-to-end latency presented in Fig. 8, as batch size increases, the batch latency of both QRMark and Stable Signature grows, but the growth rate of QRMark is significantly slower. At a batch size of 512, the latency of Stable Signature approaches nearly four times that of QRMark, highlighting the superior scalability of QRMark. The improvements in speed and latency are largely attributable to three key algorithmic and system-level optimizations: image tiling, adaptive multi-channel horizontal fusion, and resource-aware scheduling. Image tiling reduces both computational and memory costs of decoding, as each tile in our setting is only $1/16$ the size of the original image. To further optimize execution, we employ adaptive multi-channel horizontal fusion with resource-aware scheduling, which distributes CUDA streams across pipeline stages in proportion to their measured demands and assigns minibatches for each stream. This ensures that the most time-consuming stages receive more parallel capacity. The resulting multi-stream execution improves GPU utilization, boosts throughput, and lowers latency. Besides, the workload interleaving strategy, kernel fusion (full details in Appendix A.1), and optimizations on RS Correction also contribute to the improvement.
Figure 7: End-to-end image throughput comparison for (a) Stable Signature and (b) AquaLoRA with and without our QRMark under varying image batch sizes on a NVIDIA H100 GPU.
Figure 8: End-to-end watermark extraction latency comparison for (a) Stable Signature and (b) AquaLoRA with and without our QRMark under varying image batch sizes on a NVIDIA H100 GPU.
@@
Moreover, as shown in Fig. 7 and  8, QRMark further optimized AquaLoRA’s throughput and latency by fusing the image preprocessing stages into a single CUDA kernel and reducing the intermediate global-memory writes and reads between kernels. QRMark with a batch size of 32 achieves a $2.0\times$ speedup over the original AquaLoRA baseline and a $1.2\times$ speedup when using the same batch size. Notably, when scaling the batch size above 128, we observed that the original AquaLoRA baseline encounters CUDA out-of-memory errors, whereas QRMark continues to sustain high throughput and reduce latency. This further highlights our method’s advantage in large-batch image watermark decoding tasks.
@@ 8.2 Optimization Analysis
Throughput Improvement Breakdown: To better illustrate the benefits of individual optimizations and quantify their contributions to end-to-end performance, we present a detailed throughput breakdown of QRMark in Fig. 9. All experiments use a fixed message length of 48 bits and, when applicable, a tile size of $64$. Large-batch processing (LB) alone provides a modest $1.06\times$ improvement, while tiling (T) without LB yields a clearer benefit of $1.18\times$. Combining T and LB further increases the gain to $1.23\times$. Beyond these algorithmic optimizations, the system-level techniques deliver the dominant share of improvements: incorporating optimized Reed–Solomon decoding (RS) raises the speedup to $1.57\times$, and the addition of horizontal fusion and scheduling (FS) ultimately achieves a $2.43\times$ overall throughput improvement. These results show that although tiling and batching offer meaningful acceleration, the system-level RS and AS optimizations are the key factors driving high performance.
Figure 9: Throughput breakdown and incremental improvement from algorithmic and system-level optimizations. LB: large-batch processing; T: tiling; RS: optimized Reed-Solomon decoding; FS: horizontal fusion and scheduling.
Overhead Analysis: Our profiling results show that end-to-end runtime is overwhelmingly dominated by the decoding stage, which accounts for 79.0% of total execution time. This confirms that the convolutional kernels in the extractor form the primary computational bottleneck. In comparison, preprocessing and tiling contribute 14.4% and 6.2%, respectively, indicating that image normalization and tile partitioning incur relatively small overheads. Synchronization and data-transfer costs constitute only 0.4%, demonstrating that cross-stage coordination is not a limiting factor for pipeline throughput. This low synchronization cost aligns with our system design: during Reed–Solomon correction, each decoding worker transfers only a compact 48-bit message, keeping communication overhead minimal. Overall, the breakdown indicates that efficiency is constrained mainly by compute-bound decoding rather than synchronization, motivating our focus on GPU-centric and pipeline-level optimizations.
Figure 10: Scalability comparison of QRMark under data parallelism and data parallelism with horizontal fusion enabled.
Scalability Analysis: To highlight the scalability of QRMark under multi-GPU settings, we report throughput across different parallel configurations using a tile size of 64. As shown in Fig. 10, scaling with standard data parallelism exhibits inherently weak scaling for this workload. Unlike traditional DL training, the per-request computation in watermark detection is dominated by a fixed extractor that runs independently on each GPU, leaving little opportunity for cross-device parallel overlap. As a result, the baseline speedup plateaus at only $1.6\times$, $2.8\times$, and $3.3\times$ with 2, 4, and 8 GPUs, respectively. QRMark’s data-parallel design with multi-channel horizontal fusion restructures the execution pipeline to expose additional concurrency. By overlapping decoding, tiling, and Reed–Solomon correction across dedicated CUDA streams and process-level workers, the system reduces idle periods on each device and increases effective parallelism. Consequently, QRMark achieves stronger scaling, reaching $2.1\times$, $3.6\times$, and $4.1\times$ speedup at the same GPU counts—up to $30\%$ higher than the standard DP baseline. These results demonstrate that coordinated stream-level scheduling, rather than replicating the full pipeline across GPUs, is essential for converting additional hardware into meaningful throughput gains.
9 Related Work
Image Watermarking: Image watermarking is a fundamental technique for protecting and authenticating intellectual property in image synthesis and distribution. Existing approaches can be broadly categorized into three classes: post-generation, pre-generation, and in-generation watermarking [hu2024stable]. Post-generation methods [rahman2013dwt, tancik2020stegastamp, zhu2018hidden, luo2020distortion, al2007combined, jing2021hinet, zhang2020udh, zhang2019robust] embed watermarks after image creation, for instance using DWT–DCT variants, GAN-based hiding, or encoder–decoder schemes. Pre-generation watermarking embeds watermarks into all training images [yu2021artificial, zhao2023recipe], but this strategy incurs high computational cost when applied to large-scale diffusion models. In-generation watermarking [fernandez2023stable, zhao2023recipe, xiong2023flexible, feng2024aqualora, wen2023tree] integrates watermark embedding into the generation process itself—for example, Stable Signature modifies the VAE [kingma2013auto] stage, while AquaLorRA and Tree-Ring embed messages directly during diffusion sampling. Owing to its strong robustness and relatively high detection efficiency, we focus on in-generation watermarking. QRMark builds on Stable Signature and AquaLoRA, using their original implementations as baselines to demonstrate our efficiency improvements.
Reed–Solomon Error Correction: Reed–Solomon (RS) codes, introduced by Reed and Solomon in 1960 [reed1960polynomial], are linear block codes over finite fields with the maximum-distance-separable (MDS) property. They are widely deployed in both consumer and mission-critical technologies, including barcodes, QR Codes, telecommunications, and broadcasting, where they ensure reliable decoding even in the presence of noise or partial corruption. In watermarking applications, RS codes strengthen robustness by encoding watermark payloads with redundancy, enabling recovery even when some image tiles are degraded by cropping, color modifications, or noise [abdul2013error, liu2025robust]. In QRMark, RS codes are employed to mitigate accuracy loss introduced by tiling.
10 Conclusion
Appendix A Optimization and design in QRMark
A.1 Kernel Fusion for Preprocessing
A.2 ML-based Tile Size Predictor
In §4, we illustrate the pipeline for image watermark detection. However, in real‑world scenarios, a significant challenge arises: given a target image, how can we determine its watermark’s tile size? A naive strategy might involve sliding tiles of varying sizes over the image and processing each through multiple decoders pretrained for specific tile sizes. Such an approach, however, incurs prohibitive computational overhead and is impractical for real‑time deployment. To address this challenge, we propose a ML‑based Watermark Tile‑Size Predictor that efficiently estimates the watermark tile size from a given image in a single forward pass. Our predictor combines convolutional neural network feature extraction with gradient‑boosted regression, selecting an EfficientNet [tan2019efficientnet] + XGBoost [chen2016xgboost] architecture for its balance of performance and interoperability. Concretely, we adopt a pretrained EfficientNet as a backbone to extract a compact feature vector from each input image. These feature embeddings are then fed into an XGBoost regressor, which predicts the tile size with high accuracy. Similar to our preprocessing latency predictor, both training‑data collection and model training are conducted offline, thereby eliminating runtime profiling overhead.
A.3 Ablation Study on QRMark Robustness
Table 3: Validation accuracy w/ different tiling strategies.
Tiling Strategy 16 32 48 64 80
Random 0.93 0.99 0.99 0.99 0.99
Random Grid 0.75 0.97 0.99 0.99 0.99
Fixed 0.81 0.84 0.85 0.99 0.99
Table 4: Validation accuracy under different watermark #bits.
#bits 40 48 56 64 72 80 96
Bit Acc. 0.99 0.99 0.98 0.91 0.89 0.84 0.77
Word Acc. 0.99 0.99 0.29 0.00 0.00 0.00 0.00